# Optuna vs Code Llama: Are LLMs a New Paradigm for Hyperparameter Tuning?

### v3

@@ Abstract
Optimal hyperparameter selection is critical for maximizing neural network performance, especially as models grow in complexity. This work investigates the viability of leveraging large language models (LLMs) for hyperparameter optimization by fine-tuning a parameter-efficient version of Code Llama using LoRA. The adapted LLM is capable of generating accurate and efficient hyperparameter recommendations tailored to diverse neural network architectures. Unlike traditional approaches such as Optuna, which rely on computationally intensive trial-and-error procedures, our method achieves competitive or superior results in terms of Root Mean Square Error (RMSE) while significantly reducing computational overhead. Our findings demonstrate that LLM-based optimization not only matches the performance of state-of-the-art techniques like Tree-structured Parzen Estimators (TPE) but also substantially accelerates the tuning process. This positions LLMs as a promising alternative for rapid experimentation, particularly in resource-constrained environments such as edge devices and mobile platforms, where computational efficiency is essential. In addition to improved efficiency, the method offers time savings and consistent performance across various tasks, highlighting its robustness and generalizability. All generated hyperparameters are included in the LEMUR Neural Network (NN) Dataset, which is publicly available and serves as an open-source benchmark for hyperparameter optimization research.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated the ability to capture complex patterns across diverse tasks. As their scale and sophistication continue to grow, they are increasingly supplanting traditional methods in numerous domains [18, 40, 26]. With fine-tuning techniques, LLMs can acquire new capabilities and solve previously unseen problems. In this work, we explore fine-tuning LLMs for hyperparameter optimization (HPO), positioning them as a novel alternative to traditional methods such as Bayesian Optimization (BO) and Evolutionary Strategies (ES).
Our approach leverages a fine-tuned version of Code Llama to recommend hyperparameters for deep learning models in a one-shot manner. We focus on training hyperparameters for well-known computer vision architectures in the context of image classification, aiming to evaluate whether LLMs can match or surpass the performance of traditional optimization frameworks such as Optuna.
@@ 1.1 Related work
Early model-free approaches to the hyperparameter optimization (HPO) task, such as grid and random search, perform well in low-dimensional hyperparameter spaces. However, as the number of hyperparameters increases, the computational cost of applying these methods also rises. For further insights into HPO, we refer readers to [17, 36, 4, 6, 5, 9].
More advanced techniques for HPO include evolutionary strategies (ES) and Bayesian optimization (BO). Although ES outperforms earlier methods in more complex search spaces, its application can still be computationally expensive [9, 10].
BO has proven to be an efficient alternative, especially in high-dimensional hyperparameter spaces, and typically surpasses ES in terms of computational efficiency. It operates by constructing a surrogate model, updating its posterior distribution based on observed data, and selecting the next candidate using an acquisition function. The effectiveness of BO strongly depends on the surrogate model used, with the tree-structured Parzen Estimator (TPE) being the focus of this study [7]. TPE models the objective function by creating probability distributions for both favorable and unfavorable hyperparameter values. These distributions are iteratively refined during optimization, guiding the search toward regions with a higher likelihood of success. For a more detailed discussion of BO, we refer to [41, 17, 42, 8, 15, 25].
Applying this framework, Optuna [2] is a flexible hyperparameter optimization tool that supports various surrogate models within the BO paradigm. For this study, we compare Optuna’s TPE-based approach—due to its effectiveness in navigating complex hyperparameter spaces—with the Python version of Code Llama (7B and 13B parameter versions) [39].
@@
Extensive research has been conducted to evaluate whether LLMs can outperform established HPO methods. For instance, [52] demonstrated that LLMs can achieve performance comparable to or better than conventional methods when applied to models such as logistic regression, support vector machines, random forests, and neural networks, even when those methods are paired with various optimization strategies (e.g., [1]). Their study compared various HPO algorithms, including BO with Gaussian process and random forest surrogates, to a random search baseline. [12] investigated using LLMs to tune the step-size of the (1+1)-ES algorithm, showing that LLM-driven strategies can effectively compete with traditional methods. [30] introduced AgentHPO, an iterative HPO tool powered by LLMs, which frequently outperforms human-crafted solutions while maintaining interpretability. [31] proposed LLAMBO, a method that integrates LLMs with BO, demonstrating effective zero-shot warm-starting by leveraging LLMs’ domain knowledge. Lastly, [33] presented Sequential Large Language Model-Based Optimization (SLLMBO), showcasing how LLMs can outperform traditional BO methods in several scenarios.
The distinguishing feature of our study is the application of fine-tuning to the LLM for the HPO task. Unlike previous work such as [33], we focus on locally deployable models. While we acknowledge that our setup has resource limitations, to the best of our knowledge, the findings and methodology presented here are novel and have not been previously explored.
@@ 2.1 Problem Formulation
Hyperparameter tuning aims to find an optimal configuration $\lambda^{*}$ within the search space $\Lambda$ that minimizes the loss:
@@ 2.3 Fine-Tuning Process
We enhance Code Llama’s [38] hyperparameter prediction capability via LoRA [21] fine-tuning, which modifies only a subset of parameters to minimize overhead. The optimization objective is:
@@ 2.4 Evaluation Strategy
We compare our fine-tuned LLM to Optuna by measuring the RMSE of accuracy-based errors. Specifically, for each trial $i$, we define $\epsilon_{i}=1-a_{i}$, where $a_{i}$ is the observed accuracy. The RMSE is then computed as $\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\epsilon_{i})^{2}})$, providing a measure of how closely the trials match ideal (error-free) performance. Additionally, we assess stability across trials and analyze the impact of fine-tuning cycles. The goal is to demonstrate that Code Llama provides competitive hyperparameter recommendations with reduced computational cost.
@@ 3 Implementation
In this section, we elaborate on the process we implemented to yield our results. This process consisted mainly of 8 main parts depicted in Figure 1. Each of these parts is explained in its respective sub-section below.
@@ 3.1 Dataset Preparation and Initial Hyperparameter Tuning
Additionally, we prepared the source code for the following neural network models: RNN, Long Short-Term Memory (LSTM), and LLaMA 3, targeting text generation tasks. These models were trained on the Salesforce/WikiText dataset taken from the HuggingFace platform.
To determine the optimal hyperparameters of each model, we employed the Optuna framework [2]. The hyperparameters being tuned included the learning rate, batch size, and momentum, each sampled from specific ranges or sets of values. The learning rate was chosen from a continuous range between 0.0001 and 1, while the batch size was selected from the set $\{4,5,8,16,32,64\}$. The momentum, which influences the convergence speed of the training process, was selected from a range of 0.01 to 0.99. Each model was trained using these hyperparameter configurations over different numbers of epochs, namely 1, 2, and 5 (for some models) for each configuration. After the completion of each training run, the selected hyperparameter values and the accuracy achieved were recorded. The results, including the final accuracy for each set of hyperparameters, were then saved into a JSON file, providing a structured dataset of hyperparameter configurations and their corresponding performance metrics. This data collection process enabled a thorough analysis of how different hyperparameter settings influenced the performance of both image classification and text generation models.
This dataset preparation resulted in 3700 entries including all the hyperparameter-accuracy pairs determined by Optuna for each of the 17 models mentioned.
The resulting dataset was released publicly as part of the LEMUR NN Dataset [19], which serves as an open-source benchmark for research in hyperparameter optimization and AutoML.
@@ 3.2 Code Llama And Fine-Tuning
For this project, the Code-Llama-Python [39] model was chosen as the base model for fine-tuning to recommend optimal hyperparameters for neural network architectures. This decision was guided by several key factors, including open-source availability, performance and the fact that this version of Code Llama is predominantly trained on Python code.
To generate valid and meaningful hyperparameter suggestions using Code-Llama-Python, we fine-tuned the base model starting with its pre-trained checkpoint available on HuggingFace. For fine-tuning, we employed the LoRA technique [21], using a rank of 32, alpha set to 16, and a dropout rate of 0.05. The training was conducted over 35 epochs.
@@
Using our dataset we created, as explained in Section 3.1, a prompt format to use for the fine-tuning including a system prompt telling the model its role as a hyperparameter suggestion tool, an instruction asking for the three hyperparameters we are tuning for a given model mentioning its implementation code to achieve a mentioned accuracy and a response which includes the hyperparameters that would achieve the named accuracy in the instruction. The prompt format we chose is the following common choice for fine-tuning shown in Listing LABEL:listing:prompt_format.
After each cycle of fine-tuning, the LLM’s performance is evaluated by prompting it to suggest the best possible training hyperparameters for given models. The existing dataset is expanded by adding hyperparameter values generated by the Code-Llama-Python model. After the generation of these values, each set of hyperparameters undergoes validation by running a training process for the specific model at hand. The generated hyperparameters, consisting of learning rate, momentum, and batch size, are applied to the training configuration. The resulting model accuracy is then compared against the initially desired accuracy specified in the dataset. This comparison serves as a feedback loop, where the obtained accuracy is used to refine and further train the Code-Llama-Python model, improving its hyperparameter prediction capabilities for subsequent iterations. This process ensures that the model evolves to produce increasingly optimized hyperparameter sets, leading to enhanced performance in future predictions.
@@ 4.1 Pre-Trained Code Llama
Upon analyzing the responses from the original pre-trained version of Code Llama, we observed that the model returned a total of 1006 relevant records out of 2980 responses, yielding a relevance rate of approximately 33.77%. This indicates that roughly one-third of the responses were deemed pertinent to the queries posed. However, a notable observation was the presence of a significant number of answers that lacked numerical values, which could impact the utility of the responses for tasks requiring quantitative analysis.
Additionally, we identified certain patterns in the parameters suggested by the pre-trained Code Llama. For instance, the frequently suggested batch size value was 0, which raises concerns about its applicability in real-world scenarios. Typically, a batch size of 0 is impractical and may lead to confusion among users attempting to implement the model’s recommendations. Furthermore, the most common learning rate suggested in the responses was 0.01. This learning rate is often a standard starting point in many machine learning contexts, but without fine-tuning, its effectiveness may vary depending on the specific dataset and task at hand.
@@ 4.2 Baseline Results and Training Efficiency
While accuracy is essential for our tasks, RMSE was chosen for its ability to capture fine-grained error dynamics. RMSE provides a continuous measure of performance variability across trials and fine-tuning cycles, offering insights into model stability and robustness.
Fine-tuned Code Llama consistently achieves lower RMSE values than the baseline Optuna results for most models used in fine-tuning, demonstrating its effectiveness in hyperparameter optimization. As shown in Figure 2, models such as SwinTransformer, SqueezeNet, and VGG experience significant error reductions after fine-tuning, highlighting substantial improvements in hyperparameter recommendations. Additionally, architectures like ConvNeXt and GoogLeNet show moderate yet meaningful enhancements, further validating Code Llama’s adaptability across different network types. The contraction of shaded 95% confidence intervals for most architectures indicates improved stability and reduced variability in hyperparameter predictions. Since no substantial RMSE reductions were observed beyond the second fine-tuning cycle, further fine-tuning was deemed unnecessary.
@@
The results in Table 1 show that fine-tuning using Code Llama resulted in an improvement in RMSE compared to the Optuna baseline. After the first round of fine-tuning, the RMSE decreased from 0.589 to 0.563, and the confidence interval narrowed, indicating increased stability. The second round of fine-tuning was able to maintain these improvements with minor changes, while the standard deviation also decreased, confirming the increased stability of the model. In this context, the “one-shot” prediction refers to the initial result obtained from the fine-tuned Code Llama after completing its first fine-tuning cycle for a specific model. The largest reduction in RMSE was observed in the one-shot mode (to 0.533), demonstrating the potential of this approach. Further details on the one-shot approach can be found in Section 4.4.
@@
The 95% confidence intervals in this study were computed using Student’s t-distribution. Specifically, the standard deviation ($\sigma$) of errors was calculated for each trial, and the standard error ($SE=\sigma/\sqrt{n}$, where $n$ is the number of trials) was used to determine the interval bounds. The final confidence intervals were obtained as $RMSE\pm t_{\alpha/2,n-1}\times SE$, where $t_{\alpha/2,n-1}$ is the critical value from the t-distribution. This method provides a statistically rigorous estimate of uncertainty around the reported RMSE values. For more information on confidence intervals in statistical context refer to [13].

### v4

@@ Abstract
Optimal hyperparameter selection is critical for maximizing the performance of neural networks in computer vision, particularly as architectures become more complex. This work explores the use of large language models (LLMs) for hyperparameter optimization by fine-tuning a parameter-efficient version of Code Llama using LoRA. The resulting model produces accurate and computationally efficient hyperparameter recommendations across a wide range of vision architectures. Unlike traditional methods such as Optuna, which rely on resource-intensive trial-and-error procedures, our approach achieves competitive or superior Root Mean Square Error (RMSE) while substantially reducing computational overhead. Importantly, the models evaluated span image-centric tasks such as classification, detection, and segmentation, fundamental components in many image manipulation pipelines including enhancement, restoration, and style transfer. Our results demonstrate that LLM-based optimization not only rivals established Bayesian methods like Tree-structured Parzen Estimators (TPE), but also accelerates tuning for real-world applications requiring perceptual quality and low-latency processing. All generated configurations are publicly available in the LEMUR Neural Network Dataset (https://github.com/ABrain-One/nn-dataset), which serves as an open source benchmark for hyperparameter optimization research and provides a practical resource to improve training efficiency in image manipulation systems.
@@ 1 Introduction
Large Language Models (LLMs) [13] have recently demonstrated the ability to capture complex patterns across various tasks. As their scale and sophistication continue to grow, they are increasingly replacing traditional methods in numerous domains [23, 56, 37]. With fine-tuning techniques, LLMs can acquire new capabilities and solve previously unseen problems. In this work, we explore fine-tuning LLMs for hyperparameter optimization (HPO) [22], positioning them as a novel alternative to traditional methods such as Bayesian Optimization (BO) [11] and Evolutionary Strategies (ES) [25].
Our approach leverages a fine-tuned version of Code Llama [55] to recommend hyperparameters for deep learning models in a one-shot manner. We focus on training hyperparameters for well-known computer vision architectures in the context of image classification, aiming to evaluate whether LLMs can match or surpass the performance of traditional optimization frameworks such as Optuna [2].
Importantly, many of the evaluated architectures, such as ResNet [27], MobileNet [29], and SwinTransformer [45], are widely used as backbones in image manipulation pipelines, including tasks like super-resolution, deblurring, and perceptual enhancement. Efficient hyperparameter tuning of these models has a direct impact on the performance and usability of image restoration systems, particularly in resource-constrained or real-time scenarios. Thus, our findings contribute not only to HPO research, but also to the broader image manipulation community.
@@ 1.1 Related work
Early model-free approaches to the HPO task, such as grid and random search, perform well in low-dimensional hyperparameter spaces. However, as the number of hyperparameters increases, the computational cost of applying these methods also increases. For further information on HPO, we refer to [22, 51, 4, 6, 5, 9].
More advanced techniques for HPO include evolutionary strategies (ES) and Bayesian optimization (BO). Although ES outperforms previous methods in more complex search spaces, its application can still be computationally expensive [9, 10].
BO has proven to be an efficient alternative, especially in high-dimensional hyperparameter spaces, and typically exceeds ES in terms of computational efficiency. It operates by constructing a surrogate model, updating its posterior distribution based on observed data, and selecting the next candidate using an acquisition function. The effectiveness of BO strongly depends on the surrogate model used, with the tree-structured Parzen Estimator (TPE) being the focus of this study [7]. TPE models the objective function by creating probability distributions for both favorable and unfavorable hyperparameter values. These distributions are iteratively refined during optimization, guiding the search toward regions with a higher likelihood of success. For a more detailed discussion of BO, we refer to [58, 22, 60, 8, 20, 36].
Applying this framework, Optuna is a flexible hyperparameter optimization tool that supports various surrogate models within the BO paradigm. For this study, we compare Optuna’s TPE-based approach – due to its effectiveness in navigating complex hyperparameter spaces – with the Python version of Code Llama (7B and 13B parameter versions).
@@
Extensive research has been conducted to evaluate whether LLMs can outperform established HPO methods. For instance, [73] demonstrated that LLMs can achieve performance comparable to or better than conventional methods when applied to models such as logistic regression, support vector machines, random forests, and neural networks, even when those methods are paired with various optimization strategies (e.g., [1]). Their study compared various HPO algorithms, including BO with Gaussian process and random forest surrogates, to a random search baseline. [14] investigated using LLMs to tune the step-size of the (1+1)-ES algorithm, showing that LLM-driven strategies can effectively compete with traditional methods. [42] introduced AgentHPO, an iterative LLM-powered HPO tool, which frequently outperforms human-crafted solutions while maintaining interpretability. [43] proposed LLAMBO, a method that integrates LLMs with BO, demonstrating effective zero-shot warm-starting by leveraging LLMs’ domain knowledge. Lastly, [47] presented Sequential Large Language Model-Based Optimization (SLLMBO), showcasing how LLMs can outperform traditional BO methods in several scenarios.
The distinguishing feature of our study is the application of fine-tuning to the LLM for the HPO task. Unlike previous work such as [47], we focus on locally deployable models. Although we acknowledge that our setup has resource limitations, to the best of our knowledge, the findings and methodology presented here are novel and have not been previously explored.
Unlike most prior LLM-based HPO studies, our work explores the integration of such methods directly into computer vision pipelines relevant for image restoration and enhancement. We emphasize compatibility with tasks commonly found in AIM workflows, such as super-resolution and denoising, where rapid tuning of hyperparameters can substantially impact perceptual quality.
@@ 2.1 Problem Formulation
The hyperparameter tuning aims to find an optimal configuration $\lambda^{*}$ within the search space $\Lambda$ that minimizes the loss:
@@ 2.3 Fine-Tuning Process
We enhance Code Llama’s [54] hyperparameter prediction capability through LoRA [30] fine-tuning, which modifies only a subset of parameters to minimize overhead. The optimization objective is the following:
@@ 2.4 Evaluation Strategy
We compare our fine-tuned LLM with Optuna by measuring the RMSE of accuracy-based errors. Specifically, for each trial $i$, we define $\epsilon_{i}=1-a_{i}$, where $a_{i}$ is the observed accuracy. The RMSE is then computed as $\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\epsilon_{i})^{2}})$, providing a measure of how closely the trials match ideal (error-free) performance. In addition, we assess stability across trials and analyze the impact of fine-tuning cycles. The goal is to demonstrate that Code Llama provides competitive hyperparameter recommendations with reduced computational cost.
@@ 3 Implementation
In this section, we elaborate on the process we implemented to obtain our results. This process consisted mainly of eight main parts depicted in Figure 1. Each of these parts is explained in its respective subsection below.
@@ 3.1 Dataset Preparation and Initial Hyperparameter Tuning
Additionally, we prepared the source code for the following neural network models: RNN [18], Long Short-Term Memory (LSTM) [28], and Llama 3 [19], targeting text generation tasks. These models were trained on the Salesforce/WikiText dataset [49] taken from the HuggingFace platform.
To determine the optimal hyperparameters of each model, we used the Optuna framework [2]. The hyperparameters being tuned included the learning rate, batch size, and momentum, each sampled from specific ranges or sets of values. The learning rate was chosen from a continuous range between 0.0001 and 1, while the batch size was selected from the set $\{4,5,8,16,32,64\}$. The momentum, which influences the convergence speed of the training process, was selected from a range of 0.01 to 0.99. Each model was trained using these hyperparameter configurations over different numbers of epochs, namely 1, 2, and 5 (for some models) for each configuration. After completion of each training run, the selected hyperparameter values and the accuracy achieved were recorded. The results, including the final accuracy for each set of hyperparameters, were then saved in a JSON file, providing a structured dataset of hyperparameter configurations and their corresponding performance metrics. This data collection process enabled a thorough analysis of how different hyperparameter settings influenced the performance of both image classification and text generation models.
This dataset preparation resulted in 3700 entries including all the hyperparameter-accuracy pairs determined by Optuna for each of the 17 models mentioned. The resulting dataset was released publicly as part of the LEMUR NN Dataset [24], which serves as an open source benchmark for research in hyperparameter optimization and AutoML.
@@ 3.2 Code Llama And Fine-Tuning
For this project, the Code-Llama-Python [55] model was chosen as the base model for fine-tuning to recommend optimal hyperparameters for neural network architectures. This decision was guided by several key factors, including open-source availability, performance, and the fact that this version of Code Llama is predominantly trained on Python code.
To generate valid and meaningful hyperparameter suggestions using Code-Llama-Python, we fine-tuned the base model starting with its pre-trained checkpoint available on HuggingFace. For fine-tuning, we employed the LoRA technique [30], using a rank of 32, an alpha set to 16, and a dropout rate of 0.05. The training was conducted over 35 epochs.
@@
Using our dataset we created, as explained in Section LABEL:dataset-prep, a prompt format to use for the fine-tuning including a system prompt telling the model its role as a hyperparameter suggestion tool, an instruction asking for the three hyperparameters we are tuning for a given model mentioning its implementation code to achieve a mentioned accuracy and a response which includes the hyperparameters that would achieve the named accuracy in the instruction. The prompt format we chose is the following common choice for fine-tuning shown in Listing LABEL:listing:prompt_format.
After each fine-tuning cycle, the LLM’s performance is evaluated by prompting it to suggest the best possible training hyperparameters for the given models. The existing dataset is expanded by adding hyperparameter values generated by the Code-Llama-Python model. After the generation of these values, each set of hyperparameters undergoes validation by running a training process for the specific model at hand. The hyperparameters generated, consisting of learning rate, momentum, and batch size, are applied to the training configuration. The resulting model accuracy is then compared with the initially desired accuracy specified in the dataset. This comparison serves as a feedback loop, where the accuracy obtained is used to refine and further train the Code-Llama-Python model, improving its hyperparameter prediction capabilities for subsequent iterations. This process ensures that the model evolves to produce increasingly optimized hyperparameter sets, leading to enhanced performance in future predictions.
@@ 4.1 Pre-Trained Code Llama
Upon analyzing the responses from the original pre-trained version of Code Llama, we observed that the model returned a total of 1006 relevant records out of 2980 responses, yielding a relevance rate of approximately 33.77%. This indicates that approximately one-third of the responses were deemed pertinent to the queries posed. However, a notable observation was the presence of a significant number of answers that lacked numerical values, which could affect the utility of the responses for tasks that require quantitative analysis.
Additionally, we identified certain patterns in the parameters suggested by the pre-trained Code Llama. For instance, the frequently suggested batch size value was 0, raising concerns about its applicability in real-world scenarios. Typically, a batch size of 0 is impractical and can cause confusion among users who attempt to implement the recommendations of the model. Furthermore, the most common learning rate suggested in the responses was 0.01. This learning rate is often a standard starting point in many machine learning contexts, but without fine-tuning, its effectiveness may vary depending on the specific dataset and task at hand.
@@ 4.2 Baseline Results and Training Efficiency
Although accuracy is essential for our tasks, RMSE was chosen for its ability to capture fine-grained error dynamics. RMSE provides a continuous measure of performance variability across trials and fine-tuning cycles, providing insights into model stability and robustness.
The fine-tuned Code Llama consistently achieves lower RMSE values than the baseline Optuna results for most models used in fine-tuning, demonstrating its effectiveness in hyperparameter optimization. As shown in Figure 2, models such as SwinTransformer [45], SqueezeNet [32], and VGG [59] experience significant error reductions after fine-tuning, highlighting substantial improvements in hyperparameter recommendations. Furthermore, architectures such as ConvNeXt [46] and GoogLeNet [61] show moderate but meaningful enhancements, further validating the adaptability of Code Llama across different types of networks. The contraction of shaded 95% confidence intervals for most architectures indicates improved stability and reduced variability in hyperparameter predictions. Since no substantial reductions in RMSE were observed beyond the second fine-tuning cycle, further fine-tuning was deemed unnecessary.
@@
The results in Table 1 show that fine-tuning using Code Llama resulted in an improvement in RMSE compared to the Optuna baseline. After the first round of fine-tuning, the RMSE decreased from 0.589 to 0.563, and the confidence interval narrowed, indicating increased stability. The second round of fine-tuning was able to maintain these improvements with minor changes, while the standard deviation also decreased, confirming the increased stability of the model. In this context, the “one-shot” prediction refers to the initial result obtained from the fine-tuned Code Llama after completing its first fine-tuning cycle for a specific model. The largest reduction in RMSE was observed in the one-shot mode (to 0.533), demonstrating the potential of this approach. More details on the one-shot approach can be found in Section 4.4.
@@
The 95% confidence intervals in this study were calculated using Student’s t-distribution. Specifically, the standard deviation ($\sigma$) of the errors was calculated for each trial and the standard error ($SE=\sigma/\sqrt{n}$, where $n$ is the number of trials) was used to determine the interval bounds. The final confidence intervals were obtained as $RMSE\pm t_{\alpha/2,n-1}\times SE$, where $t_{\alpha/2,n-1}$ is the critical value from the t-distribution. This method provides a statistically rigorous estimate of the uncertainty around the reported RMSE values. For more information on confidence intervals in statistical context, refer to [15].